# Keep it Light! Simplifying Image Clustering via Text-Free Adapters

### v1

@@ 3.1 Pair Construction Backbone
The success of BYOL demonstrates that we can maximize the similarities of positive pairs without negative ones. In SCP, the positive pairs consist of samples augmented from the same instance.
Given a data instance $x_{i}$, we apply two stochastic transformations $T^{a}$ and $T^{b}$, independently selected from the augmentation family $\mathcal{T}$. This produces two correlated views: $x_{i}^{a}=T^{a}(x_{i})$ and $x_{i}^{b}=T^{b}(x_{i})$.
An appropriate augmentation strategy is vital for better downstream performance. In our work, we adopt only two simple augmentations: RandomCrop and GaussianBlur. This choice aligns with the preprocessing techniques used in training pre-trained models, ensuring compatibility with their learned representations. RandomCrop randomly crops the image to a specified size, and GaussianBlur applies a Gaussian filter to blur the image. For each image, these two augmentations are applied independently, each with a 50% probability. We then use a pre-trained model $f(\cdot)$, such as CLIP, to extract features from the augmented images: $h_{i}^{a}=f(x_{i}^{a})$ and $h_{i}^{b}=f(x_{i}^{b}).$
@@ 3.2 Cluster Head
Following the “label as representation” concept [23], when a data sample is projected into a space whose dimensionality matches the number of clusters $K$, the $k$-th component of its feature vector (after applying a $\operatorname{softmax}$ function) can be interpreted as the probability that the sample belongs to the $k$-th cluster. We employ a five-layer non-linear MLP as the clustering head $g(\cdot)$, producing a $K$-dimensional feature that is normalized with a $\operatorname{softmax}$ over the dimension of the cluster.
@@
Hence, $y_{i}^{a}$ and $y_{i}^{b}$ are both $K$-dimensional vectors, whose components $y_{i,k}^{a}$ and $y_{i,k}^{b}$ indicate the probability of assigning the $i$-th sample to the $k$-th cluster. Formally, let $Y^{a},Y^{b}\in\mathbb{R}^{N\times K}$ be the outputs of the clustering head for all samples. Then, we have the following matrices:
$Y^{a}=\begin{bmatrix}y^{i}_{1}\\ \vdots\\ y^{a}_{N}\end{bmatrix}\quad Y^{b}=\begin{bmatrix}y^{b}_{1}\\ \vdots\\ y^{b}_{N}\end{bmatrix}.$
To maximize row-wise similarity, we adopt the following cross-entropy loss function instead of the commonly used InfoNCE loss [20], as SCP only have positive pairs that should share similar soft assignments:
@@
Inspired by the effective regularizations in TAC [18], we further introduce the following confidence loss to make the soft labels $y^{a}_{i}$ and $y^{b}_{i}$ more confident, approaching one-hot vectors:
@@
In addition, following TAC [18], we introduce an entropy term $H(Y)$ to prevent model collapse, defined as follows:
@@
where
$P^{a}_{k}=\frac{1}{N}\sum_{i=1}^{N}y^{a}_{i,k},\quad P^{b}_{k}=\frac{1}{N}\sum_{i=1}^{N}y^{b}_{i,k}.$
$H(Y)$ encourages uniform soft assignments across clusters, thereby mitigating the issue of empty clusters.
Hence, we define the overall objective function of SCP as
@@
where the balancing weight $\alpha$ modulates the influence of $H(Y)$, especially when the number of clusters is large. By maximizing consistency between different augmented views with regularizations, SCP effectively prevents trivial solutions and achieves competitive performance. We provide algorithm 1 to explain our pipeline.
@@
1:Dataset $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$, Pre-trained model $f(\cdot)$, number of clusters $K$, batch size $B$, loss weight $\alpha$
2:Initialize cluster head $g(\cdot)$
3:for each epoch do
4:     for each mini-batch $\{x_{i}\}_{i=1}^{B}$ do
5:         Pair Construction:
6:         for each data instance $x_{i}$ in the mini-batch do
7:              Apply stochastic transformations $T^{a}$, $T^{b}$ to obtain:
8:               $x_{i}^{a}=T^{a}(x_{i})$,  $x_{i}^{b}=T^{b}(x_{i})$
9:              Extract features using pre-trained model:
10:               $h_{i}^{a}=f(x_{i}^{a})$,  $h_{i}^{b}=f(x_{i}^{b})$
11:         end for
12:         Cluster Space Encoding:
13:         for each feature $h_{i}^{a}$, $h_{i}^{b}$ do
14:              Compute soft assignments:  $y_{i}^{a}=g(h_{i}^{a})$,  $y_{i}^{b}=g(h_{i}^{b})$
15:         end for
16:         Compute Losses:
17:         Compute total clustering loss:
18:          $L_{\text{clu}}=L_{e}+L_{\text{con}}-\alpha H(Y)$
19:         Update cluster head $g(\cdot)$ parameters by minimizing $L_{\text{clu}}$
20:     end for
21:end for
22:return soft assignments $y_{i}=g(h_{i})$ for each $x_{i}\in\mathcal{X}$
@@ 4 Experiments
In this section, we evaluate the proposed SCP on six widely challenging image clustering datasets. A series of initial quantitative and qualitative comparisons, ablation studies, and hyper-parameter analyses will be carried out to investigate the effectiveness and robustness of the method.
@@ 4.1.1 Datasets
To evaluate the performance of SCP, we apply it to six widely used image clustering datasets: STL-10 [38], CIFAR-10 [26], CIFAR-20 [26], CIFAR-100 [26], ImageNet-10 [39], and ImageNet-Dogs [39], which is a subset from ImageNet-1k [40]. The brief information of all datasets used in our evaluation is summarized in Table 1.
Table 1: Characteristics of the benchmark datasets used in our evaluation.
@@ 4.1.2 Implementation Details
In most experiments, we utilized the CLIP [15] with the backbone ViT-B/32 [30]. For a fair comparison with TEMI [36] and CPP [35], we replaced the CLIP backbone with ViT-L/14 to match the architectures used in these methods. The cluster head $g$ is a five-layer MLP with dimensions: $D\rightarrow 1024\rightarrow 786\rightarrow 512\rightarrow 1024\rightarrow K$, where $D$ denotes the output dimension of the pre-trained model and $K$ represents the number of clusters.
@@ 4.2 Main Results
We compare our method with state-of-the-art baselines on six widely-used image clustering datasets, with feature visualizations to show the competitiveness of our proposed SCP.
4.2.1 Comparison against classical and text-based methods
We first evaluate our pipeline on five widely used datasets and compare it with 18 deep clustering baselines. Since most earlier baselines adopt ResNet-34 or ResNet-18 as the backbone, we mainly focus on comparisons with CLIP and CLIP-based methods. Specifically, CLIP (zero-shot) uses CLIP’s pre-trained image and text encoders to classify images by matching them to text prompts, while CLIP (K-means) clusters directly on CLIP image features via K-means. Our approach includes SCP-CLIP, which employs a ViT-B/32 backbone like other CLIP-based methods, and SCP-DINO, which uses a ViT-B/8 backbone. Each experiment is repeated 50 times, and the best results are reported.
Table 2: Clustering performance of various baseline methods on multiple datasets (all metrics are multiplied by 100). We highlight the best and second-best results in boldfaced red and underlined in blue, respectively. CLIP (K-means), CLIP (zero-shot), SCP-CLIP, SIC, and TAC all adopt a ViT-B/32 backbone. SCP-DINO uses a Vit-B/8 backbone.
@@ 4.2.1 Comparison against classical and text-based methods
JULE [41] 18.2 27.7 16.4 19.2 27.2 13.8 10.3 13.7 3.3 17.5 30.0 13.8 5.4 13.8 2.8
DEC [42] 27.6 35.9 18.6 25.7 30.1 16.1 13.6 18.5 5.0 28.2 38.1 20.3 12.2 19.5 7.9
DAC [39] 36.6 47.0 25.7 39.6 52.2 30.6 18.5 23.8 8.8 39.4 52.7 30.2 21.9 27.5 11.1
DCCM [43] 37.6 48.2 26.2 49.6 62.3 40.8 28.5 32.7 17.3 60.8 71.0 55.5 32.1 38.3 18.2
IIC [44] 49.6 59.6 39.7 51.3 61.7 41.1 22.5 25.7 11.7 - - - - - -
PICA [45] 61.1 71.3 53.1 59.1 69.6 51.2 31.0 33.7 17.1 80.2 87.0 76.1 35.2 35.3 20.1
CC [23] 76.4 85.0 72.6 70.5 79.0 63.7 43.1 42.9 26.6 85.9 89.3 82.2 44.5 42.9 27.4
IDFD [46] 64.3 75.6 57.5 71.1 81.5 66.3 42.6 42.5 26.4 89.8 95.4 90.1 54.6 59.1 41.3
SCAN [10] 69.8 80.9 64.6 79.7 88.3 77.2 48.6 50.7 33.3 - - - 61.2 59.3 45.7
MiCE [47] 63.5 75.2 57.5 73.7 83.5 69.8 43.6 44.0 28.0 - - - 42.3 43.9 28.6
GCC [48] 68.4 78.8 63.1 76.4 85.6 72.8 47.2 47.2 30.5 84.2 90.1 82.2 49.0 52.6 36.2
NNM [22] 66.3 76.8 59.6 73.7 83.7 69.4 48.0 45.9 30.2 - - - 60.4 58.6 44.9
TCC [24] 73.2 81.4 68.9 79.0 90.6 73.3 47.9 49.1 31.2 84.8 89.7 82.5 55.4 59.5 41.7
SPICE [25] 81.7 90.8 81.2 73.4 83.8 70.5 44.8 46.8 29.4 82.8 92.1 83.6 57.2 64.6 47.9
SIC [19] 95.3 98.1 95.9 84.7 92.6 84.4 59.3 58.3 43.9 97.0 98.2 96.1 69.0 69.7 55.8
TAC [18] 95.5 98.2 96.1 83.3 91.9 83.1 61.1 60.7 44.8 98.5 99.2 98.3 80.6 83.0 72.2
SCP-CLIP 94.6 97.9 95.1 85.5 93.2 85.7 59.5 60.1 44.3 97.4 98.5 97.9 56.3 59.6 44.6
@@
(a) Step 1, $ARI=4.5$
(b) Step 10, $ARI=24$
(c) Step 50, $ARI=57$
(d) Step 100, $ARI=59$
Figure 2: Visualization of representations from different training steps learned by SCP-DINO on the ImageNet-Dogs training set, along with the corresponding clustering ARI (multiplied by 100). (a) Image embeddings directly from the DINO image encoder, with clusters obtained using K-means. (b)–(d) Image logits learned by SCP-DINO across different steps. There are $77$ steps per epoch.
As shown in Table 2, our method significantly enhances clustering performance across multiple benchmark datasets. For example, on CIFAR-10, CLIP (K-means) achieves an ACC of 75.2%, whereas SCP-CLIP improves this by 18%, surpassing TAC by 1.3%. On CIFAR-20, SCP-CLIP attains an ACC of 60.1%, outperforming SIC by 1.8%. Moreover, SCP-DINO boosts performance on four benchmarks, achieving state-of-the-art results on STL-10 and the second-best performance on both ImageNet-10 and ImageNet-Dogs. The lower performance of SCP-CLIP on ImageNet-Dogs may stem from the difficulty of capturing discriminative features in fine-grained images, particularly without aligned text features.
Notably, SCP relies exclusively on visual features but still surpasses CLIP zero-shot on all benchmarks and outperforms TAC on both STL-10 and CIFAR-10. Results underscore its effectiveness in extracting and leveraging visual information for clustering. Furthermore, our pipeline is text-free, featuring a simpler architecture with fewer hyperparameters—making it especially advantageous for adaptation to pure visual pre-trained models.
Overall, these results demonstrate that beyond zero-shot classification or multi-modal frameworks, a simpler approach can still effectively utilize the power of pre-trained models for image clustering.
@@ 4.2.2 Comparison with recent text-free methods
To supplement our answers to Question (Q1), we introduce SCP-MIX, which leverages the concatenated visual features from both CLIP and DINO. Alongside SCP-CLIP, we compare this method against two recent text-free approaches, TEMI [36] and CPP [35], as reported in Table 3.
Table 3: Clustering performance comparison for recent text-free methods. The best and second-best results are highlighted in boldface red and underlined in blue, respectively. All methods use the ViT-L/14 backbone, and SCP-MIX adopts the concatenation (or ensemble) of features learned by CLIP (ViT-L/14) and DINO (ViT-B/8).
@@
SCP-MIX outperforms CPP and TEMI on both CIFAR-10 and CIFAR-20, while SCP-CLIP achieves the highest accuracy on CIFAR-100. These results underscore the competitive performance of SCP, highlighting effectiveness and simplicity. Notably, SCP achieves these outcomes without relying on additional feature layers, support sets, self-distillation, or exponential moving average strategies. Thus, unimodal text-free deep clustering method can rival multimodal pipelines.

### v2

@@ 3.1 Pair Construction Backbone
The success of BYOL demonstrates that we can maximize the similarities of positive pairs without negative ones. In SCP, the positive pairs consist of samples augmented from the same instance. Given a data instance $x_{i}$, we apply two stochastic transformations $T^{a}$ and $T^{b}$, independently selected from the augmentation family $\mathcal{T}$. This produces two correlated views: $x_{i}^{a}=T^{a}(x_{i})$ and $x_{i}^{b}=T^{b}(x_{i})$.
Although comprehensive augmentation strategies could enhance downstream performance, we adopt only two simple and commonly used augmentations in our work: RandomCrop and GaussianBlur. This choice aligns with the preprocessing techniques used in training pre-trained models, ensuring compatibility with their learned representations. RandomCrop randomly crops the image to a specified size, and GaussianBlur applies a Gaussian filter to blur the image. For each image, these two augmentations are applied independently, each with a 50% probability. We then use a pre-trained model $f(\cdot)$, such as CLIP, to extract features from the augmented images: $h_{i}^{a}=f(x_{i}^{a})$ and $h_{i}^{b}=f(x_{i}^{b}).$
@@ 3.2 Cluster Head
Following the “label as representation” concept [li2021contrastive], when a data sample is projected into a space whose dimensionality matches the number of clusters $K$, the $k$-th component of its feature vector (after applying a $\operatorname{softmax}$ function) can be interpreted as the probability that the sample belongs to the $k$-th cluster. We employ a five-layer non-linear MLP as the clustering head $g(\cdot)$, producing a $K$-dimensional feature that is normalized with a $\operatorname{softmax}$ over the dimension of the cluster.
@@
Hence, $y_{i}^{a}$ and $y_{i}^{b}$ are both $K$-dimensional vectors, whose components $y_{i,k}^{a}$ and $y_{i,k}^{b}$ indicate the probability of assigning the $i$-th sample to the $k$-th cluster.
To maximize row-wise similarity, we adopt the following cross-entropy loss function instead of the commonly used InfoNCE loss [oord2018representation], as SCP only have positive pairs that should share similar soft assignments:
@@
Inspired by the effective regularizations in previous works such as TAC [li2023image], we further introduce the following confidence loss to make the soft labels $y^{a}_{i}$ and $y^{b}_{i}$ more confident, approaching one-hot vectors:
@@
In addition, following TAC [li2023image], we introduce an entropy term $H(Y)$ to prevent model collapse, defined as follows:
@@
where $P^{a}_{k}=\frac{1}{N}\sum_{i=1}^{N}y^{a}_{i,k},\quad P^{b}_{k}=\frac{1}{N}\sum_{i=1}^{N}y^{b}_{i,k}.$ And $H(Y)$ encourages uniform soft assignments across clusters, thereby mitigating the issue of empty clusters. Hence, we define the overall objective function of SCP as
@@
where the balancing weight $\alpha$ modulates the influence of $H(Y)$, especially when the number of clusters is large. By maximizing consistency between different augmented views with regularizations, SCP effectively prevents trivial solutions and achieves competitive performance. We also provide Algorithm 1 to explain our pipeline.
Input: $\mathcal{X}=\{x_{i}\}_{i=1}^{N}$, pre-trained model $f(\cdot)$, clusters $K$, batch size $B$, loss weight $\alpha$
Output: Soft assignments $y_{i}=g(h_{i})$ for each $x_{i}$
Initialize cluster head $g(\cdot)$;
for each epoch do
 for each mini-batch $\{x_{i}\}_{i=1}^{B}$ do
    /* Pair Construction */
    for each $x_{i}$ in mini-batch do
       $x_{i}^{a}=T^{a}(x_{i})$, $x_{i}^{b}=T^{b}(x_{i})$;
       $h_{i}^{a}=f(x_{i}^{a})$, $h_{i}^{b}=f(x_{i}^{b})$;
    /* Cluster Space Encoding */
    for each $h_{i}^{a},h_{i}^{b}$ do
       $y_{i}^{a}=g(h_{i}^{a})$, $y_{i}^{b}=g(h_{i}^{b})$;
    /* Compute Losses */
    $L_{\mathrm{clu}}=L_{e}+L_{\mathrm{con}}-\alpha H(Y)$;
      Update parameters of $g(\cdot)$ by minimizing $L_{\mathrm{clu}}$;
@@ 4 Experiments
In this section, we evaluate the proposed SCP on eight widely used image clustering datasets. A series of initial quantitative and qualitative comparisons, ablation studies, and hyper-parameter analyses will be carried out to investigate the effectiveness and robustness of the method.
@@ 4.1.1 Datasets
To evaluate the performance of SCP, we apply it to six widely used image clustering datasets: STL-10 [coates2011analysis], CIFAR-10 [krizhevsky2009learning], CIFAR-20 [krizhevsky2009learning], CIFAR-100 [krizhevsky2009learning], ImageNet-10 [chang2017deep], and ImageNet-Dogs [chang2017deep], which is a subset of ImageNet-1k [deng2009imagenet]. In addition, we include two more challenging datasets, DTD [cimpoi2013describingtextureswild] and UCF-101 [soomro2012ucf101dataset101human]. The brief information of all datasets used in our main evaluation is summarized in the following table:
TABLE II: Characteristics of the benchmark datasets used in our evaluation.
@@
DTD Train+Val/Test 3,760 1,880 47
UCF-101 Train/Val 9,537 3,783 101
4.1.2 Evaluation Metrics
We use three widely adopted metrics Accuracy (ACC), Normalized Mutual Information (NMI), and Adjusted Rand Index (ARI) to evaluate clustering performance. Higher values of these metrics indicate better results.
@@ 4.1.3 Implementation Details
In most experiments, we used the original OpenAI CLIP [radford2021learning] with the ViT-B/32 backbone [dosovitskiy2020image]. For a fair comparison with TEMI [adaloglou2023exploring] and CPP [chu2024image], we replaced the CLIP backbone with ViT-L/14 to match the architectures used in those methods. SCP-DINO employs the ViT-B/8 architecture from original DINO [caron2021emerging], while SCP-MIX adopts concatenated features from both CLIP and DINO. The clustering head $g$ is a five-layer MLP with the following architecture: $D\rightarrow 1024\rightarrow 786\rightarrow 512\rightarrow 1024\rightarrow K$, where $D$ is the output dimension of the pre-trained model and $K$ is the number of clusters.
@@ 4.2 Main Results
We compare our method with state-of-the-art baselines on six widely used image clustering datasets and provide feature visualizations to demonstrate that competitive performance can be achieved with a simple deep clustering pipeline.
4.2.1 Comparison against classical and text-based method
We first evaluate our pipeline on five widely used datasets and compare it against 18 deep clustering baselines. Since most earlier baselines adopt ResNet-34 or ResNet-18 as their backbone, we primarily focus on comparisons with CLIP and CLIP-based methods. Specifically, CLIP (zero-shot) uses CLIP’s pre-trained image and text encoders to classify images by matching them to text prompts, while CLIP (K-means) performs clustering directly on CLIP image features using K-means. Our approach includes SCP-CLIP, which employs a ViT-B/32 backbone, consistent with other CLIP-based methods. To support our claim that SOTA clustering performance can be achieved using sufficiently informative image representations alone—and to demonstrate the flexibility of SCP—we also include SCP-DINO, which adopts the ViT-B/8 backbone. In addition, SCP-MIX further combines features from both DINO and CLIP via simple concatenation.
TABLE III: Clustering performance of various baseline methods on multiple datasets (all metrics are multiplied by 100). We highlight the best and second-best results in bold and underlined, respectively. CLIP (K-means), CLIP (zero-shot), SCP-CLIP, SIC, and TAC all adopt a ViT-B/32 backbone. SCP-DINO uses a ViT-B/8 backbone. SCP-MIX adopts the concatenation of features learned by CLIP (ViT-B/32) and DINO (ViT-B/8).
@@ 4.2.1 Comparison against classical and text-based method
JULE [yang2016jointunsupervisedlearningdeep] 18.2 27.7 16.4 19.2 27.2 13.8 10.3 13.7 3.3 17.5 30.0 13.8 5.4 13.8 2.8
DEC [xie2016unsuperviseddeepembeddingclustering] 27.6 35.9 18.6 25.7 30.1 16.1 13.6 18.5 5.0 28.2 38.1 20.3 12.2 19.5 7.9
DAC [chang2017deep] 36.6 47.0 25.7 39.6 52.2 30.6 18.5 23.8 8.8 39.4 52.7 30.2 21.9 27.5 11.1
DCCM [wu2019deepcomprehensivecorrelationmining] 37.6 48.2 26.2 49.6 62.3 40.8 28.5 32.7 17.3 60.8 71.0 55.5 32.1 38.3 18.2
IIC [ji2019invariantinformationclusteringunsupervised] 49.6 59.6 39.7 51.3 61.7 41.1 22.5 25.7 11.7 - - - - - -
PICA [huang2020deep] 61.1 71.3 53.1 59.1 69.6 51.2 31.0 33.7 17.1 80.2 87.0 76.1 35.2 35.3 20.1
CC [li2021contrastive] 76.4 85.0 72.6 70.5 79.0 63.7 43.1 42.9 26.6 85.9 89.3 82.2 44.5 42.9 27.4
IDFD [tao2021clusteringfriendlyrepresentationlearninginstance] 64.3 75.6 57.5 71.1 81.5 66.3 42.6 42.5 26.4 89.8 95.4 90.1 54.6 59.1 41.3
SCAN [van2020scan] 69.8 80.9 64.6 79.7 88.3 77.2 48.6 50.7 33.3 - - - 61.2 59.3 45.7
MiCE [tsai2020mice] 63.5 75.2 57.5 73.7 83.5 69.8 43.6 44.0 28.0 - - - 42.3 43.9 28.6
GCC [zhong2021graph] 68.4 78.8 63.1 76.4 85.6 72.8 47.2 47.2 30.5 84.2 90.1 82.2 49.0 52.6 36.2
NNM [dang2021nearest] 66.3 76.8 59.6 73.7 83.7 69.4 48.0 45.9 30.2 - - - 60.4 58.6 44.9
TCC [shen2021you] 73.2 81.4 68.9 79.0 90.6 73.3 47.9 49.1 31.2 84.8 89.7 82.5 55.4 59.5 41.7
SPICE [niu2022spice] 81.7 90.8 81.2 73.4 83.8 70.5 44.8 46.8 29.4 82.8 92.1 83.6 57.2 64.6 47.9
SIC [cai2023semantic] 95.3 98.1 95.9 84.7 92.6 84.4 59.3 58.3 43.9 97.0 98.2 96.1 69.0 69.7 55.8
TAC [li2023image] 95.5 98.2 96.1 83.3 91.9 83.1 61.1 60.7 44.8 98.5 99.2 98.3 80.6 83.0 72.2
SCP-CLIP 94.6 97.9 95.1 85.5 93.3 85.7 59.5 60.1 44.3 97.4 98.5 97.9 56.3 59.6 44.6
@@
SCP-MIX 97.9 98.9 97.6 91.7 96.5 92.5 62.7 60.6 45.3 98.0 99.3 98.5 80.6 84.0 74.5
As shown in Table III, our method significantly enhances clustering performance across multiple benchmark datasets. For example, on CIFAR-10, CLIP (K-means) achieves an ACC of 75.2%, whereas SCP-CLIP improves this by 18%, surpassing TAC by 1.3%. On CIFAR-20, SCP-CLIP attains an ACC of 60.1%, outperforming SIC by 1.8%. The relatively lower performance of SCP-CLIP on ImageNet-Dogs may stem from the difficulty of capturing discriminative features in fine-grained images, particularly in the absence of aligned textual information. However, SCP-DINO, which takes advantage of a powerful vision backbone, mitigates this limitation and achieves an ACC of 80.5%. With a slight increase in complexity, SCP-MIX further boosts performance across all benchmarks, achieving state-of-the-art accuracy on STL-10, CIFAR-10, ImageNet-10, and ImageNet-Dogs—demonstrating the potential of SCP’s structure and further supporting our claim.
Notably, SCP-CLIP relies exclusively on visual features yet still surpasses CLIP zero-shot on the most benchmarks, and achieves competitive performance even when compared to TAC on STL-10, CIFAR-10, and CIFAR-20. These results underscore its effectiveness in extracting and leveraging visual representations for clustering. Furthermore, our pipeline is entirely text-free, featuring a simpler architecture with fewer hyperparameters—making it especially advantageous for adaptation to purely visual pre-trained models.
Overall, these results demonstrate that beyond zero-shot classification and multimodal frameworks, a simple and lightweight approach can still achieve competitive clustering performance. Moreover, SCP is flexible and applicable to a variety of pre-trained models. To explore its applicability in more diverse settings, we evaluate SCP on two additional challenging datasets and three recent pre-trained models.
@@ 4.2.2 Comparison with recent text-free methods
TABLE IV: Clustering performance comparison for recent text-free methods. The best and second-best results are highlighted in bold and underlined, respectively. All methods use the ViT-L/14 backbone, and SCP-MIX adopts the concatenation of features learned by CLIP (ViT-L/14) and DINO (ViT-B/8).
@@
Compared to recent unimodal competitors, under the same settings, SCP-CLIP achieves highly competitive performance across all CIFAR benchmarks with a simpler design. With additional representations, SCP-MIX further improves the results and outperforms methods such as CPP and TEMI on both CIFAR-10 and CIFAR-20. These results underscore the effectiveness and simplicity of the SCP framework. In particular, SCP achieves these results without relying on additional feature layers, support sets, self-distillation, or exponential moving average strategies during training. This simple yet effective design makes SCP accessible to the deep clustering community—serving as both a strong text-free competitor and a practical alternative when text-guided methods are not applicable.
Performance on Challenging Datasets
In addition to benchmark datasets, we also tested SCP on more challenging datasets that reflect more complicated image types. UCF-101 [soomro2012ucf101dataset101human] is a video dataset focused on human actions rather than distinct objects, reflecting more dynamic and complex visual content. We extract one frame from each clip and split the data into training and validation sets. DTD [cimpoi2013describingtextureswild] contains texture-centric images without outstanding central objects, offering a different type of visual structure. We set $\alpha=2$ for DTD and $\alpha=3$ for UCF-101; other settings remain the same.
TABLE V: Clustering performance on DTD and UCF-101. Best and second-best are in bold and underline, respectively. SCP-CLIP uses ViT-B/32; SCP-DINO uses ViT-B/8. SCP-MIX concatenates CLIP (ViT-B/32) and DINO (ViT-B/8) features.
Method DTD (NMI / ARI / ACC) UCF-101 (NMI / ARI / ACC)
SIC [cai2023semantic] 59.6 / 45.9 / 30.5 81.0 / 61.9 / 53.6
TAC [li2023image] 62.1 / 50.1 / 34.4 82.3 / 68.7 / 60.1
SCP-CLIP 58.4 / 46.1 / 30.3 77.7 / 61.2 / 51.4
SCP-DINO 57.5 / 45.3 / 29.6 80.9 / 64.0 / 54.4
SCP-MIX 61.3 / 50.2 / 34.3 81.3 / 66.3 / 56.5
CLIP (K-means) 57.3 / 42.6 / 27.4 79.5 / 58.2 / 47.6
CLIP (zero-shot) 56.5 / 43.1 / 26.9 79.9 / 63.4 / 50.2
We observe that SCP still achieves competitive results on both DTD and UCF-101 without text guidance, showing its capability for broader image clustering tasks when text is unavailable.
Applicability of SCP
To further demonstrate the applicability of SCP, we compare its performance using three additional backbones—SigLIP v2 [tschannen2025siglip2multilingualvisionlanguage], BLIP-2 [li2023blip2bootstrappinglanguageimagepretraining], and MAE [he2021maskedautoencodersscalablevision]—across all benchmark and challenging datasets in Table  VI . We observe that SCP consistently improves clustering performance in most cases compared to the K-means baselines.
TABLE VI: Clustering performance on multiple datasets compared with different backbones. We compare the K-means baseline with SCP-enhanced variant. Arrows indicate improvement () or degradation () over the K-means baseline.
Method STL-10 CIFAR-10 CIFAR-20 ImageNet-10 ImageNet-Dogs DTD UCF-101
NMI ARI ACC NMI ARI ACC NMI ARI ACC NMI ARI ACC NMI ARI ACC NMI ARI ACC NMI ARI ACC
SigLIPv2 (K-means) 92.0 84.5 83.2 82.5 79.0 70.7 60.6 52.3 37.0 91.1 79.8 84.7 75.6 73.1 59.4 64.2 50.9 37.0 76.3 46.4 34.4
SCP-SigLIPv2 96.2 98.2 96.2 92.6 96.8 93.2 64.5 62.4 48.4 98.2 99.3 98.5 82.4 85.4 76.8 65.4 53.1 38.8 80.0 67.7 59.0
BLIP-2 (K-means) 86.6 76.8 73.8 93.3 84.7 85.3 75.1 67.7 57.2 99.1 99.7 99.4 73.5 67.4 57.5 65.5 52.3 37.4 84.8 74.0 66.6
SCP-BLIP-2 97.7 98.4 96.6 97.8 99.2 98.3 74.5 72.0 60.7 99.3 99.8 99.5 85.2 88.7 81.6 64.8 52.7 38.6 88.4 75.5 68.9
MAE (K-means) 44.2 40.2 30.2 40.8 43.6 27.8 31.6 26.4 15.6 63.5 69.6 50.5 3.6 11.8 1.5 27.4 12.7 6.1 47.1 26.7 14.7
SCP-MAE 53.9 59.7 40.7 48.6 58.4 39.4 34.4 35.0 19.5 84.7 91.7 83.4 11.0 16.2 4.3 38.9 22.9 10.3 47.6 26.1 15.1
Particularly, in real-world applications, SCP could easily be equipped and fit to an advanced backbone or a local representation learning approach to optimize the clustering due to the simple design.